OpenStudy (anonymous):

What is 4.4% is in a mixed faction or faction in simplest form?

OpenStudy (skullpatrol):

any ideas?

OpenStudy (anonymous):

It's not a fraction at all.

OpenStudy (anonymous):

In a frasction it would be 4.4 over a 100.

OpenStudy (anonymous):

I need get the percent into fraction or mixed one.

OpenStudy (anonymous):

I lied

OpenStudy (anonymous):

4.4 = 4.4/1 4.4/1 * 10/10 = 44/10 44/10 ==> 22/5

OpenStudy (skullpatrol):

$$\Huge \ 4.4 \text{ percent} = \frac{4.4}{100} $$

OpenStudy (skullpatrol):

percent means divide by 100

OpenStudy (skullpatrol):

Ok so far?

OpenStudy (skullpatrol):

$$percent = *\frac{1}{100}$$